# Randomization Tests for Conditional Group Symmetry

### v1

@@ 6 Kernel hypothesis tests
Our CRT implementation, shown in Algorithm 3, is able to use either of these statistics, or the adaptive MMD-based statistics described in the next section. The basic idea is to simulate conditionally independent randomizations $Z_{1:n}^{(b)}$ and ${Z^{\prime}}_{1:n}^{(b)}$, then use them to estimate the MMD. This produces a null distribution with which to compare the estimated MMD between $Z_{1:n}$ and yet another independent randomization ${Z}^{(0)}_{1:n}$. Optionally, a single set of comparison samples $Z^{\prime}_{1:n}$ can be reused (option reuse in Algorithm 3) instead of independent comparison samples ${Z}_{1:n}^{(b)}$, $b=0,\dotsc,B$. The level of the test in each case is $\alpha$, though we expect different power properties; we study both versions in the experiments in Section 7.
In the theoretical analysis of the MMD-based tests, not all of the usual statistical properties of MMD-based tests hold. Whereas in the standard setting independent samples are compared, here the samples are dependent through conditioning on $S_{1:n}$ for the randomization. The two samples are, however, conditionally independent, which allows us to build on established properties of MMD estimators when independent samples are used. For ease of theoretical analysis of the MMD-based test, we assume that independent comparison sets (reuse = false) and the V-statistic are used. We use the U-statistic in our experiments, but we do not expect the two to behave very differently.
@@
$\Delta_{n}^{2}\coloneq\mathbb{E}\left[\widehat{\mathrm{MMD}}_{v,\sigma}^{2}(Z_{1:{n}},{Z}_{1:{n}}^{(0)})\;\bigg{|}\;S_{1:n}\right]$ (18)
$=\frac{1}{n^{2}}\sum_{i,j}\mathbb{E}\left[k_{\sigma}(Z_{i},Z_{j})+k_{\sigma}(Z^{(0)}_{i},Z^{(0)}_{j})-2k_{\sigma}(Z_{i},Z^{(0)}_{j})\mid S_{i},S_{j}\right]\;.$
If $\Delta_{n}$ is sufficiently large, we expect the test to have power against non-equivariant alternatives. To state the formal result, let $\nu^{2}=\sup_{z}k_{\sigma}(z,z)<\infty$, and define for $\eta>0$,
$R_{n}(\eta)\coloneq\frac{9}{2n}\nu^{2}+\frac{16\nu^{2}\sqrt{2}}{\sqrt{n^{3}}}+6\nu\sqrt{\frac{2}{n}}+\frac{3\nu}{2}\sqrt{\eta}\sqrt{\frac{2}{n}+32\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta\nu}{n}\;.$ (19)
Because we analyze the conditional (on $S_{1:n}$) test, probabilistic properties of the test can be said to hold conditionally (on $S_{1:n}$), almost surely (with respect to $S_{1:n}$). Analysis of bootstrap procedures rely on the same ideas (e.g., van der Vaart, 1998, Ch. 23).
@@
Let $T$ be $\widehat{\mathrm{MMD}}_{v,\sigma}^{2}$, with kernel $k_{\sigma}$ and corresponding RKHS $\mathcal{H}_{k_{\sigma}}$, using exact randomization in the CRT. Assume that $\Delta_{n}\geq 2\nu\sqrt{2/n}$. Then for every $\eta>0$, for almost every sequence $S_{1:n}$, conditionally on $S_{1:n}$, with probability at least $1-e^{-\eta}$ over observations $Z_{1:n}$ and randomization ${Z}^{(0)}_{1:n}$,
@@
$p_{\eta}(\Delta_{n})=\exp\left(-\frac{3}{2}\Delta_{n}+R_{n}(\eta)\right)\;.$ (21)
The proof can be found in Section B.4. Observe that for fixed $\alpha$, $B$, and $n$, the lower bound in (20) is increasing in $\Delta_{n}$. Because multiple approximations were used for the sake of deriving explicit bounds, tighter lower bounds seem possible for (21) and in Corollary 6.2 below, but we do not pursue them here.
Using the Chernoff bound on the upper tail of deviations of a binomial random variable, we can obtain an explicit lower bound on values of $\Delta_{n}$ that correspond to alternatives for which the test (with high probability) rejects the null with probability at least $\beta\in(0,1)$. For convenience, let $B_{\alpha}\coloneq\lfloor\alpha(B+1)-1\rfloor$.
Corollary 6.2.
$\Delta_{n}\geq\frac{2}{3}\left(R_{n}(\eta)-\ln\frac{B_{\alpha}}{B}-\ln\left(1-\sqrt{\frac{-2\ln(1-\beta)}{B_{\alpha}}}\right)\right)\;,$ (22)
then for almost every sequence $(S_{i})_{i\geq 1}$, conditionally on $S_{1:n}$, we have that $\mathbb{P}(p_{B}\leq\alpha|S_{1:n})\geq\beta$ with probability over $(Z_{1:n},{Z}^{(0)}_{1:n})$ at least $1-e^{-\eta}$.
In the limit of $B,n\to\infty$, we see that if $\text{MMD}(P,P^{(0)})\geq-\frac{2}{3}\ln\alpha$, then the test will reject with probability $\beta$ for any $\beta\in(0,1)$.
@@ 6.1 Adaptive kernel test statistics
It is known that the power of kernel-based tests are highly dependent on the chosen bandwidth parameter $\sigma$ (Ramdas et al., 2015). Several solutions have been proposed, in the form of data-based heuristics (Gretton et al., 2012; Garreau, Jitkrittum and Kanagawa, 2017) and, more recently, adaptive methods (Schrab et al., 2023; Biggs, Schrab and Gretton, 2024; Cárcamo, Cuevas and Rodríguez, 2024). In our experiments, when using a kernel with a parameter (e.g., the Gaussian kernel), we use two test statistics based on solutions proposed in the adaptive kernel literature. The first adapts the unnormalized FUSE statistic
@@
where in practice $\omega=\sqrt{n(n-1)}$ following Biggs, Schrab and Gretton (2024). The statistic can be interpreted as taking a softmax across estimated MMD2 values computed using different kernels $k_{\sigma_{\ell}}$. The bandwidth values $\{\sigma_{1},\dotsc,\sigma_{L}\}$ in the original work (Biggs, Schrab and Gretton, 2024) are quantiles of the pairwise Euclidean distances in the data. In the context of a permutation test, the empirical measure of the data is a maximal invariant, so any function of it (such as quantiles) can be used in the context of a test for permutation-invariance, as in Theorem 2.2. Our adaptation, which satisfies the requirements of Proposition 5.1, is obtained by replacing the quantiles of the observation with quantiles of variables contained in the conditioning set $S_{1:n}$, such as the pairwise Euclidean distances in $\tilde{Y}_{1},\dotsc,\tilde{Y}_{n}$.
@@
Tests using this procedure remain valid due to the conditioning argument in Proposition 5.1.
Analogous versions of the power lower bound in Theorem 6.1 hold for the FUSE and SK statistics. We state a combined version here; separate proofs are given in Proposition B.7 and Proposition B.8, respectively, in Section B.5. For convenience, let $R_{n}^{\vee}(\eta)$ denote the maximum over $\ell\in[L]$ of $R_{n,\ell}(\eta)$, which is obtained by substituting $\nu_{\ell}$ for $\nu$ into (19). Define
$\bar{\Delta}_{n}\coloneq\frac{1}{L}\sum_{\ell=1}^{L}\sqrt{\mathbb{E}\left[\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z^{\prime(0)}_{1:n})\mid S_{1:n}\right]}\;,$
and
$\Delta_{n}^{\vee}=\sup_{\ell\in[L]}\Delta_{n,\ell}\coloneq\sup_{\ell\in[L]}\sqrt{\mathbb{E}\left[\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}^{2}(Z_{1:{n}},{Z}_{1:{n}}^{(0)})\mid S_{1:n}\right]}\;.$
Proposition 6.3.
Let $T$ be the FUSE (resp. SK) statistic using $\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}$ and exact randomization in the CRT. Assume that for each $\ell\in\{1,\dotsc,L\}$, $\Delta_{n,\ell}\geq 2\nu_{\ell}\sqrt{2/n}$. Then for every $\eta>0$, for almost every sequence $S_{1:n}$, conditionally on $S_{1:n}$, with probability at least $1-Le^{-\eta}$ over $(Z_{1:n},{Z}^{(0)}_{1:n})$,
$\mathbb{P}(p_{B}\leq\alpha\mid S_{1:n})\geq F_{B,p_{\eta}(\widetilde{\Delta}_{n})}(\lfloor\alpha(B+1)-1\rfloor)\;,$
$p_{\eta}(\widetilde{\Delta}_{n})=\min\left\{1,L\exp\left(-\frac{3}{2}\widetilde{\Delta}_{n}+R_{n}^{\vee}(\eta)\right)\right\}\;,$
and $\widetilde{\Delta}_{n}=\bar{\Delta}_{n}$ for FUSE, and $\widetilde{\Delta}_{n}=\Delta^{\vee}_{n}$ for SK.
Using this, a lower bound for $\widetilde{\Delta}_{n}$ that guarantees power $\beta$ with high probability, analogous to the one in Corollary 6.2, can be obtained.
@@ 7 Experiments
We evaluate our proposed tests on simple synthetic examples and on realistic simulated data from particle physics. In all of our experiments, we sample $n$ data points from a distribution or dataset and perform a test for a conditional symmetry with respect to some specified group. We repeat this procedure over $N=1000$ simulations for each test setting and report the proportion of simulations in that the test rejected, which is an estimate of the test size (resp. power) when the distribution has (resp. does not have) the specified symmetry. With $N=1000$ simulations, estimates are precise up to approximately ±$\pm 0.016$. We set the test level to be $\alpha=0.05$ in all of our experiments. Where applicable, we use $B=100$ randomizations for estimating the $p$-value in randomization tests, except when specified otherwise.
We default to aggregating Gaussian and Laplace kernels via the FUSE and SK test statistics for data that are continuous. To compute the bandwidths $\{\sigma_{1},\dotsc,\sigma_{L}\}$, we follow a procedure similar to the one described by Biggs, Schrab and Gretton (2024). We first compute the distances $\{\|\tilde{Y}_{i}-\tilde{Y}_{j}^{\prime}\|_{2}\}_{i,j}$ and $\{\|\tilde{Y}_{i}-\tilde{Y}_{j}^{\prime}\|_{1}\}_{i,j}$ for the Gaussian and Laplace kernels, respectively. We then compute half of the 5% quantile and two times the 95% quantile for each set, and take the $L=10$ bandwidth values to be the discrete uniform interpolation of those two values. All estimates of MMD use the unbiased U-statistic.
All experiments were implemented in Julia (version 1.6.1) and run on a high-performance computing allocation with 4 CPU cores and 16 GB of RAM. Additional experiments and experimental details can be found in Appendix D. Associated computer code can be found on GitHub. [^3] https://github.com/chiukenny/Tests-for-Conditional-Symmetry
@@ 7.1 Experimental configurations
In our experiments involving tests for conditional symmetry, we consider up to eight tests, each of which corresponds to a particular configuration of the following parameters:
@@
    Permutation test (BS; Section 5.2) versus conditional randomization test (RN; Section 5).
@@
    Reusing a randomized comparison set (reuse = true in Algorithm 3) versus using independent comparison sets (reuse = false).
  3. 3.
@@
Figure 6: Rejection rates for tests that reuse a comparison set in the changing covariance experiment for varying values of the data dimension $d$, $Y|X$ covariance matrix off-diagonal values $p$, and sample size $n$. The conditional distribution is $\mathrm{SO}(d)$-equivariant only for $p=0$.
@@ 7.2 Synthetic experiments
We start with several synthetic experiments involving simple distributions for the marginal of $X$ and the conditional $Y|X$. These experiments will be used to compare and demonstrate the properties of the different tests for conditional symmetry. One of these experiments is discussed in this section; the rest can be found in Appendix D.
@@
Figure 7: Rejection rates for tests that reuse a comparison set in the changing covariance experiment, where the non-$\mathrm{SO}(2)$-equivariance comes from the covariance $p$ between the first two dimensions, for varying values of the data dimension $d$ and sample size $n$.
@@
Consider random variables $X$ and $Y$ in $\mathbb{R}^{d}$. Let $X$ have a marginal distribution given by $P_{X}=\mathsf{N}(0,\Sigma_{X,d})$, where $\Sigma_{X,d}$ is the $d\times d$ matrix with $1$’s on the diagonal and $0.75$ on the off-diagonals. Let $Y|X$ have the conditional distribution $\mathsf{N}(X,\Sigma_{Y,d}(p))$, where $\Sigma_{Y,d}(p)$ is the $d\times d$ matrix with $1$’s on the diagonals and $p$ on the off-diagonals. In words, the distribution of $Y|X$ is a multivariate normal centered on $X$. For $p=0$, the distribution is spherical; as $p$ moves towards $1$, the distribution becomes more elliptical and eventually line-like. Under the action of $\mathrm{SO}(d)$, the conditional distribution is $\mathrm{SO}(d)$-equivariant if and only if $p=0$.
@@
To test for $\mathrm{SO}(d)$-equivariance on data generated from these distributions, we condition on the maximal invariant $M(X)=\|X\|$. Here, we only show results for the tests that reuse a comparison set; results for tests that use independent comparison sets in this experiment can be found in Section D.2. Figure 6 shows the estimated test rejection rates for various values of $p$ and $d$ as the sample size $n$ increases from $25$ to $500$. The randomization tests appear to be more powerful than the baseline tests in general. We also observe that increasing the dimensions leads to an increase in power, though this appears to be an artifact of the experimental setup. Consider a similar setup but where the only non-zero off-diagonal entry of $\Sigma_{Y,d}$ is between the first and second dimensions. In other words, the non-equivariance of the distribution is isolated to the first two dimensions. Figure 7 shows the rejection rates for this setting. Here, the power decreases with dimension, which suggests that the non-equivariance signal-to-noise ratio decreases with additional “equivariant dimensions”. Histograms of the $p$-values in these experiments can be found in Section D.2.
Additional synthetic experiments are found in Appendix D: under a similar setting, Section D.3 investigates the effect of the number of randomizations $B$ and Section D.4 compares sampling from an approximate conditional distribution to sampling from the true conditional; Section D.5 investigates non-equivariance in the mean as opposed to in the covariance; Section D.6 investigates tests for permutation-equivariance in another synthetic setting.

### v2

@@ 5 Kernel-based hypothesis tests for conditional group symmetry
Our CRT implementation, shown in Algorithm 3, is able to use either of these statistics or the adaptive MMD-based statistics described in the next section. The basic idea is to simulate conditionally independent randomizations $Z_{1:n}^{(b)}$ and ${Z^{\prime}}_{1:n}^{(b)}$, then use them to estimate the MMD. This produces an estimate of the test statistic under the null distribution with which to compare the estimated MMD between $Z_{1:n}$ and yet another independent randomization ${Z}^{(0)}_{1:n}$. Optionally, a single set of comparison samples $Z^{\prime}_{1:n}$ can be reused (option reuse in Algorithm 3) instead of independent comparison samples ${Z^{\prime}}_{1:n}^{(b)}$, $b=0,\dotsc,B$. The level of the test in each case is $\alpha$, though we expect different power properties.
In the theoretical analysis of the MMD-based tests, not all of the usual statistical properties of MMD-based tests hold. Whereas in the standard setting, independent samples are compared, here the samples are dependent through conditioning on $S_{1:n}$ for the randomization. The two samples are, however, conditionally independent, which allows us to build on techniques for analyzing MMD estimators when independent samples are used (Sriperumbudur et al., 2012). For ease of theoretical analysis of the MMD-based test, we assume that independent comparison sets (reuse = false) and the V-statistic are used. We use a single comparison set (reuse = true) and the U-statistic in our experiments, but we do not expect the tests to behave very differently. Simulation experiments comparing reused and independent comparison sets (see Section E.1) support this conclusion.
@@
$\Delta_{n}^{2}$ $\coloneq\mathbb{E}\left[\widehat{\mathrm{MMD}}_{v,\sigma}^{2}(Z_{1:{n}},{Z}_{1:{n}}^{(0)})\;\bigg|\;S_{1:n}\right]$
$=\frac{1}{n^{2}}\sum_{i,j}\mathbb{E}\left[k_{\sigma}(Z_{i},Z_{j})+k_{\sigma}(Z^{(0)}_{i},Z^{(0)}_{j})-2k_{\sigma}(Z_{i},Z^{(0)}_{j})\mid S_{1:n}\right]\;.$
If $\Delta_{n}$ is sufficiently large, we expect the test to have power against non-equivariant alternatives. To state the formal result, let $\kappa^{2}=\sup_{z}k_{\sigma}(z,z)<\infty$, and define for $\eta>0$,
$C_{n}(\eta)\coloneqq\kappa\left(6\sqrt{\frac{2}{n}}+3\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{n}\right)+\kappa^{2}\left(\frac{9}{n}+72\sqrt{\frac{2}{n^{3}}}\right)\;.$ (14)
Because we analyze the test that conditions on $S_{1:n}$, probabilistic properties of the test can be said to hold conditionally (on $S_{1:n}$), almost surely (with respect to $S_{1:n}$). Analysis of bootstrap procedures relies on the same ideas (e.g., van der Vaart, 1998, Ch. 23).
@@
Let $T$ be $\widehat{\mathrm{MMD}}_{v,\sigma}$, with kernel $k_{\sigma}$ and corresponding RKHS $\mathcal{H}_{k_{\sigma}}$, using exact randomization in the CRT. For any $\eta>0$, if
$\Delta_{n}>2\kappa\left(2\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\;,$ (15)
then for almost every sequence $S_{1:n}$, conditionally on $S_{1:n}$, with probability at least $1-e^{-\eta}$ over observations $Z_{1:n}$ and randomization ${Z}^{(0)}_{1:n}$,
@@
$p_{\eta}(\Delta_{n})=\min\left\{1,\exp\left(-\frac{3}{2}\Delta_{n}+C_{n}(\eta)\right)\right\}\;.$ (17)
The proof can be found in Section C.4. Observe that for fixed $\alpha$, $B$, and $n$, the lower bound in (16) is increasing in $\Delta_{n}$. Because multiple approximations were used for the sake of deriving explicit bounds, a tighter lower bound seems possible for (17), but we do not pursue that here. A lower bound on values of $\Delta_{n}$ that correspond to alternatives for which the test with high probability rejects the null with conditional probability at least $1-\beta$ can be found in Section C.4.2.
@@ 5.1 Adaptive kernel test statistics
It is known that the power of kernel MMD-based tests is highly dependent on the chosen kernel parameters, such as the bandwidth $\sigma$ (Ramdas et al., 2015). Several solutions have been proposed, in the form of data-based heuristics (Gretton et al., 2012; Garreau, Jitkrittum and Kanagawa, 2017) and, more recently, adaptive methods (Schrab et al., 2023; Biggs, Schrab and Gretton, 2024; Cárcamo, Cuevas and Rodríguez, 2024). In our experiments, when using a kernel with a parameter (e.g., the Gaussian kernel), we use two test statistics based on solutions proposed in the adaptive kernel literature. The first adapts the unnormalized FUSE statistic
@@
where, following Biggs, Schrab and Gretton (2024), $\omega=\sqrt{n(n-1)}$. The statistic can be interpreted as taking a softmax across estimated MMD2 values computed using different kernels $k_{\sigma_{\ell}}$. The bandwidth values $\{\sigma_{1},\dotsc,\sigma_{L}\}$ in the original work (Biggs, Schrab and Gretton, 2024) are quantiles of the pairwise Euclidean distances in the data. In the context of a permutation test, the empirical measure of the data is a maximal invariant, so any function of it (such as quantiles) can be used in the context of a test for permutation-invariance, as in Section 2.2. Our adaptation to CRTs for conditional group symmetry, which satisfies the requirements of Proposition 4.1, is obtained by replacing the quantiles of the observation with quantiles of variables contained in the conditioning set $S_{1:n}$, such as the pairwise Euclidean distances in $\tilde{Y}_{1},\dotsc,\tilde{Y}_{n}$.
@@
Similar to the FUSE statistic above, $\sigma_{\ell}$ can depend on $S_{1:n}$, and using this procedure remains valid due to the conditioning argument in Proposition 4.1.
Analogous versions of the power lower bound in Theorem 5.1 hold for the FUSE and SK statistics. For brevity, their statements and proofs can be found in Section C.4.
@@ 6 Experiments
We evaluate our proposed tests on synthetic examples and on realistic state-of-the-art simulated data used in particle physics. In all of our experiments, we sample $n$ data points from a distribution or dataset and perform a test for conditional symmetry with respect to a specified group. We repeat this procedure over $N=1000$ simulations for each test setting and report the proportion of simulations in which the test rejected, which is an estimate of the test size (resp. power) when the distribution has (resp. does not have) the specified conditional symmetry. With $N=1000$ simulations, estimates are precise up to approximately $\pm 0.016$. We set the test level to be $\alpha=0.05$ in all of our experiments. Except where specified otherwise, we use a single randomized comparison set (reuse = true) for estimating the $p$-value in our tests.
We default to aggregating Gaussian and Laplace kernels via the FUSE and SK test statistics for continuous data. To compute the bandwidths $\{\sigma_{1},\dotsc,\sigma_{L}\}$, we follow a procedure similar to the one described by Biggs, Schrab and Gretton (2024). We first compute the distances $\{\|\tilde{Y}_{i}-\tilde{Y}_{j}^{\prime}\|_{2}\}_{i,j}$ and $\{\|\tilde{Y}_{i}-\tilde{Y}_{j}^{\prime}\|_{1}\}_{i,j}$ for the Gaussian and Laplace kernels, respectively. We then compute half of the 5% quantile and two times the 95% quantile for each set, and take the $L=10$ bandwidth values to be the discrete uniform interpolation of those two values. All estimates of MMD use the U-statistic.
All experiments were implemented in Julia (version 1.6.1) and run on a high-performance computing allocation with 4 CPU cores and 16 GB of RAM. Additional experiments and experimental details can be found in Appendix E, including an experiment showing how our tests for conditional symmetry can be used for checking learned symmetry in predictive models (Section E.9). Associated computer code can be found on GitHub. [^4] https://github.com/chiukenny/Tests-for-Conditional-Symmetry
@@ 6.1 Experimental configurations
In our experiments involving tests for conditional symmetry, we consider up to four tests, each of which corresponds to a particular configuration of the following parameters:
@@
    Permutation test as a baseline (BS; Section 4.2) versus conditional randomization test (RN; Section 4).
@@ 6.2 Synthetic experiments
We start with several synthetic experiments involving simple distributions for the marginal of $X$ and the conditional $Y|X$. These experiments will be used to compare and demonstrate the properties of the different tests for conditional symmetry. In all experiments described in this section, we use $B=100$ randomizations for estimating the $p$-value, chosen for computational convenience due to the large number of experimental configurations evaluated. In practice, we recommend using a larger number of randomizations to minimize the risks associated with performing random inference (see Section E.3).
6.2.1 Gaussian data and rotation-equivariance
@@
Figure 4: Rejection rates for tests that reuse a comparison set in the changing covariance experiment (Section 6.2) for varying values of the data dimension $d$, $Y|X$ conditional covariance matrix off-diagonal values $c$, and sample size $n$. The conditional distribution is $\mathrm{SO}(d)$-equivariant only for $c=0$. (Note the different vertical scale for $c=0$ versus $c>0$.)
@@
Consider random variables $X$ and $Y$ in $\mathbb{R}^{d}$. Let $X$ have a marginal distribution given by $P_{X}=\mathsf{N}(0,\Sigma_{X,d})$, where $\Sigma_{X,d}$ is the $d\times d$ matrix with $1$’s on the diagonal and $0.75$ on the off-diagonals. Let $Y|X$ have the conditional distribution $\mathsf{N}(X,\Sigma_{Y,d}(c))$, where $\Sigma_{Y,d}(c)$ is the $d\times d$ matrix with $1$’s on the diagonals and $c\in[0,1)$ on the off-diagonals. In words, the distribution of $Y|X$ is a multivariate normal centered on $X$. For $c=0$, the distribution is spherical; as $c$ moves towards $1$, the distribution becomes more elliptical and eventually line-like. Under the action of $\mathrm{SO}(d)$, the conditional distribution is $\mathrm{SO}(d)$-equivariant if and only if $c=0$.
Figure 5: Rejection rates for tests that reuse a comparison set in the changing covariance experiment (Section 6.2), where the non-$\mathrm{SO}(d)$-equivariance comes from the covariance $c$ between the first two dimensions, for varying values of the data dimension $d$ and sample size $n$. (Note the different vertical scale for $c=0$ versus $c>0$.)
@@
To test for $\mathrm{SO}(d)$-equivariance on data generated from these distributions, we condition on the maximal invariant $M(X)=\|X\|$. Figure 4 shows the estimated test rejection rates for various values of $c$ and $d$ as the sample size $n$ increases from $25$ to $500$. The randomization tests appear to be more powerful than the baseline tests in general. We also observe that increasing the number of dimensions leads to an increase in power, because each dimension provides additional signal for departures from the null hypothesis. In contrast, consider a similar setup where the only non-zero off-diagonal entry of $\Sigma_{Y,d}$ is between the first and second dimensions. That is, the non-equivariance of the distribution is isolated to the first two dimensions. Figure 5 shows the rejection rates: the power decreases with dimension, which suggests that the non-equivariance signal-to-noise ratio decreases with additional “equivariant dimensions.”
6.2.2 Approximate versus exact conditional sampling
Consider a setting where $X$ is generated in the following way. Let $Z\sim\chi_{1}$ be a chi random variable with $1$ degree of freedom, and take $X\coloneq Z\Gamma\mathbf{e}_{1}$, where $\Gamma|Z,\varepsilon\sim P_{\Psi|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|Z1_{d}+\varepsilon)$ is a $d$-dimensional rotation matrix, with $1_{d}$ being the vector of ones of length $d$ and $\varepsilon\sim\mathsf{N}(0,\mathbf{I}_{d})$. Here, the conditional distribution $P_{\Gamma|R}$ is known and can be sampled from exactly. Let $Y|X$ have the same conditional distribution described in the previous section, with $c$ being the covariance between any pair of dimensions of $Y$, so that the conditional distribution of $Y|X$ is $\mathrm{SO}(d)$-equivariant if and only if $c=0$.
For $d\in\{3,4,5\}$, we compare the performance of tests that sample from an approximation of $P_{\Gamma|R}$ as per Algorithm 2 with tests that sample from $P_{\Gamma|R}$ exactly. The test results are shown in Figure 6. We see that the tests equipped with exact conditional sampling appear to have level-$\alpha$ (within reasonable variation due to sampling). Moreover, the tests with approximate sampling are comparable to those tests in terms of power, though the approximation also appears to make the tests slightly conservative. This effect is also reflected in Figure 7, which shows the estimated $p$-value distributions for these tests. In particular, in the null (equivariant) setting $(c=0)$, the estimated $p$-value distribution for the tests with exact sampling appears to be approximately uniform, and the tests with approximate sampling appear to be less uniform with slightly more mass on higher values and correspondingly lower mass on lower values. In Section E.2, we further investigate the connection between the approximation and the properties of the test; in particular, we find that Silverman’s rule for the bandwidth of the kernel used in the conditional randomization produces a conservative test in this setting, and that larger bandwidths can increase power at the risk of inflating Type I error rates.
Figure 6: Rejection rates for $\mathrm{SO}(d)$-equivariance tests for varying values of the $Y|X$ covariance matrix off-diagonal values $c$ and sample size $n$. The dotted lines represent tests that sample from the true conditional $P_{\Gamma|R}$, and the solid lines represent tests that sample from an approximation. (Note the different vertical scale for $c=0$ versus $c>0$.)
Figure 7: Histograms showing the distribution of $1000$ $p$-values of the $\mathrm{SO}(3)$-equivariant tests in the approximate versus exact sampling experiment at sample size $n=250$. Parameter $c$ is the covariance between any two dimensions.
6.2.3 Additional experiments
Additional synthetic experiments are found in Appendix E. Section E.1 compares tests that use independent comparison sets to those in Section 6.2.1 that reuse comparison sets. It finds that in addition to being less computationally expensive, reused comparison sets appear to result in a slight increase in power. Section E.3 explores the risks of using randomization-based inference using measures proposed by Stoepker and Castro (2024); there, we find that the Monte Carlo estimation of the $p$-value does not greatly interfere with the outcome of the test but does potentially result in $p$-values that are less interpretable. Section E.4 examines how using a deterministic representative inversion (if it exists) when the group action is not free can still yield a level-$\alpha$ test. Section E.5 investigates the runtimes of several variations of the FUSE test and shows that our tests are not prohibitively expensive; in particular, it suggests that setting $B$ to a value large enough to ensure accurate Monte Carlo estimates requires only a modest computational budget. Sections E.6 and E.7 consider different experimental contexts where rotation-equivariance is lost due to changes in the mean as opposed to in the covariance, and where permutation-equivariance is the symmetry of interest, respectively. In both experiments, our tests detect the alternative setting with large enough sample sizes.